OpenStudy (onepieceftw):

f(x) = sqrx+8; g(x) = 8x - 12 Find f(g(x)).

OpenStudy (maddy1251):

I can help.

OpenStudy (maddy1251):

You remember how when it was simple, we would be given a function f(x)=x+10, where we would find f(2)=2+10?

OpenStudy (anonymous):

3rd one

OpenStudy (maddy1251):

This is quite similar. For f(g(x)), our 'x' in our f(x), we will putin function g(x)

OpenStudy (anonymous):

It looks like this... 22x−1−−−−−√

OpenStudy (maddy1251):

f(x) = sqrx+8; g(x) = 8x - 12 Find f(g(x) f(x)=sqrt (x+8) = g(8x-12)= sqrt (8x-12+8)

OpenStudy (maddy1251):

@mjmahmood You're a little off.

OpenStudy (onepieceftw):

@mjmahmood Maddy is helping me and you're just trying to give me the answer. I want to learn how to solve it.

OpenStudy (maddy1251):

f(g(x)) = sqrt (8x-12+8), where now you just combine the 12 and 8 at the end.

OpenStudy (maddy1251):

Do you have answer choices to this?

OpenStudy (maddy1251):

pretty sure its f(g(x)) = sqrt (8x-4) if I combined the 12 and 8 correctly.

OpenStudy (onepieceftw):

These are my choices

OpenStudy (maddy1251):

Ah. I'm thinking A.

OpenStudy (maddy1251):

Wait

OpenStudy (maddy1251):

No, maybe C. My bad.

OpenStudy (maddy1251):

Not sure how the square root could have that power of 8. I haven't learned of that.

OpenStudy (maddy1251):

We usually always have 3 or 2, but having that 2 with answer choice c specifically should be the answer.

OpenStudy (maddy1251):

It wants you to simplify the (8x-4) because that is not in simplest form.

OpenStudy (onepieceftw):

This is precal but it's supposed to be stuff I learned a long time ago I just forgot it :/

OpenStudy (maddy1251):

Don't worry I am in pre-calc too, I am surprised at how much I have to try to remember.

OpenStudy (maddy1251):

I know it's not b. I mean, considering the way you gave me the function, it shouldn't be.

OpenStudy (maddy1251):

D doesn't make sense with the function given.

OpenStudy (maddy1251):

And C matches the actual simplified form of what would be 8x-4. Given radical properties and blah blah blah

OpenStudy (maddy1251):

Unless, my math is wrong. But I wrote it out. It's whatever you think is right. (-:

OpenStudy (onepieceftw):

I'm trying to work it out... lol Really appreciate your help :) I don't doubt you're right just need to figure out why so i can do it no problem in the future..

OpenStudy (maddy1251):

Understandable. I am the same way! It's basically finding which is the simplified form of 8x-4. I am searching stuff now to help.

OpenStudy (onepieceftw):

f(sqr(x+8)) = 8(sqr(x+8)) - 12 is what we're trying to solve, right?

OpenStudy (maddy1251):

f(x) = sqrx+8; g(x) = 8x - 12 Find f(g(x) f(8x-12)= sqrt (8x+8-12) = sqrt (8x-4) is what I get

OpenStudy (onepieceftw):

I'm totally getting backwards lol sorry about that

OpenStudy (maddy1251):

Thats okay (-:

OpenStudy (maddy1251):

I can' see or understand why it the radical would have the compendium of 8 that doesn't make sense to me, not for those functions.

OpenStudy (maddy1251):

2^3=8 so 3^sqrt(8)=2.

OpenStudy (maddy1251):

And having it be 8^sqrt would produce pretty large numbers.

OpenStudy (maddy1251):

This is what I think

OpenStudy (maddy1251):

8x-4 lets set that to f(x)

OpenStudy (maddy1251):

Then we could maybe set that to y=8x-4. Uhm. Divding it by 4 to both sides gives y=2x-1

OpenStudy (maddy1251):

f(x)= sqrt 2x-1

OpenStudy (maddy1251):

C is your answer.Just figured it out!

OpenStudy (maddy1251):

Because to simplify y=8x-4, we can put 4 into 8x and -4.

OpenStudy (maddy1251):

Or keep it f(x). I set to y because it makes any calculations easy.

OpenStudy (maddy1251):

So for this, C is your answer.

OpenStudy (onepieceftw):

Thank you :) I just to find out how it simplifies

OpenStudy (maddy1251):

Me too. I was trying to think of a way to simplify 8x-4 because it wasn't in simplest form. Then it dawned on me ' oh silly, set it to just f(x)=8x-4 and divide by 4 cause thats the common greatest factor between the two ' Which then yields 2x-1. So that only proves C to be correct, and probably after you apply 2^sqrt to 2x-1 you should get an equation similar or equal to 8x-4.

OpenStudy (onepieceftw):

I found it: since 8 is not a square we need to break it into 2*4x - 4 which simplfies to 2sqr2x-1 (the imaginary number is sqr-1 * sqr4 and I guess gets lumps with the 2 on the outside of the radical)

OpenStudy (maddy1251):

Great work there @OnePieceFTW ! :)

OpenStudy (onepieceftw):

Thank you :) Really appreciate the help!

OpenStudy (maddy1251):

Makes sense now, cause then times it by 2 gives 4 which then can be squared, etc.

OpenStudy (maddy1251):

Anytime! I think we each learned something, that's what math is about.